10) Determine the range of a so that the quadratic form Q(x, y, z) = a(r+y+z)+2xy-2yz+2zx positive definite.
Answer #1

vo) Q(X, Y, 2) = q (a+y+2>) +any - 2y2 +22x Q = XTAX where X- & to be positive definite if mateix & minors should be positiveX so => a[~~1] -2[2+1] + [-12-2]50 X 다. - 1 x²x-4-2-2-x>0 x²-24-830 (x-4) (x+2) =0 <>4072 - condemn 3 from conditions , , ③ a